# GEMeX-ThinkVG: Towards Thinking with Visual Grounding in Medical VQA via Reinforcement Learning

### v1

@@ 4.2.1 Group Relative Policy Optimization
$\quad\quad\mathcal{J}_{\text{GRPO}}(\theta)&=\mathbb{E}_{q\sim P(Q),\,\{o_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}(O|q)}\Bigg{[}\\ \frac{1}{G}\sum_{i=1}^{G}\min\Bigg{(}\frac{\pi_{\theta}(o_{i}\mid q)}{\pi_{\theta_{\text{old}}}(o_{i}\mid q)}A_{i},&\text{clip}\left(\frac{\pi_{\theta}(o_{i}\mid q)}{\pi_{\theta_{\text{old}}}(o_{i}\mid q)},1-\varepsilon,1+\varepsilon\right)A_{i}\Bigg{)}\\ &-\beta D_{\text{KL}}(\pi_{\theta}\parallel\pi_{\text{ref}})\Bigg{]}\end{aligned}$
where $A_{i}=\frac{r_{i}-\text{mean}(\{r_{1},r_{2},\ldots,r_{G}\})}{\text{std}(\{r_{1},r_{2},\ldots,r_{G}\})}$ is the relative advantage over group rewards $\{r_{i}\}_{i}^{G}$, $D_{\text{KL}}$ is the KL divergence between $\pi_{\theta}$ and $\pi_{\text{ref}}$ (the model after SFT stage) to prevent catastrophic forgetting caused by over optimization, $\epsilon$ and $\beta$ are the PPO clipping hyper-parameter [27] and the coefficient for the KL penalty, respectively.
messages = [“role”:“user”, “content”:
f“‘We would like to request your feedback on the performance of two AI assistants in response to the user question displayed above. For your reference, the visual content in the image is represented with caption describing the same image. Please rate the accuracy (most important), relevance of their responses, considering both answer and reason (if any). Each assistant receives an overall score on a scale of 1 to 10, where a higher score indicates better overall performance. Please output both scores and your reaon in JSON format { assistant1: score, assistant2: score, reason:your_reason }.”””
Table 3: Our designed prompt for Semantics Reward.
4.2.2 Reward Functions
We propose two reward functions, tailored for the two output parts of GEMeX: answer and involved visual locations. (1) Semantics Reward: In GEMeX, four types of questions are defined (Table 2). For choice-based and closed-ended questions, reward design is straightforward, as we can directly use Accuracy Reward [24] to evaluate whether the model’s outputs are correct. Nevertheless, designing a reward function for open-ended questions is particularly challenging due to the absence of fixed answers. Unlike [18], which employs traditional machine learning metrics such as BLEU as rewards—metrics that have been shown to poorly capture essential medical content [11]—we introduce a semantics-aware accuracy reward. It enables the integration of all four types into a unified, coherent reward paradigm. Specifically, we input both the ground truth (GT) answer and the model-generated answer into OpenBioLLM-70B, which can provide two accuracy scores at the semantics level. When these two scores are close, i.e., if the absolute difference between them is less than 2, we assign a reward of 1 point; otherwise, the reward is 0. The specific prompt is shown in Table 3. (2) Grounding Reward: Besides giving answers, the model needs to provide the visual locations involved in solving the questions. Therefore, we first check whether the number of bounding boxes predicted by the model matches the number in the GT. If they match, we then evaluate whether the mean Intersection over Union (mIoU) exceeds 0.75. A reward score of 1.0 is assigned only when both conditions are satisfied; otherwise, the reward is 0. Note that we leave out the common format reward as after the SFT stage, the model has learned to follow the format requirement.
5 Experiments
@@
GEMeX-ThinkVG-200K 94.71 56.52 87.11 66.23 80.08 61.76 70.20 51.13 83.03
GEMeX-ThinkVG-200K 96.44 59.63 90.97 68.39 83.31 64.52 73.18 54.49 85.98
[…]
5.1 Dataset
@@
We conduct experiments on the GEMeX dataset and randomly sample a portion of its training set to generate ThinkVG data. We denote the version with the complete training set of 1.59M VQA triples as GEMeX-Full. The 200K subset that includes ThinkVG data is denoted as GEMeX-ThinkVG-200K, while the subset without ThinkVG is referred to as GEMeX-200K. For evaluation, we use the original GEMeX test set, which consists of 300 images and 3,960 questions (1,144 open-ended questions, 543 closed-ended questions, 1,300 single-choice questions, and 973 multiple-choice questions).
5.2 Training Details
@@
We mainly explore the Qwen2.5-VL-7B-Instruct [^2] https://huggingface.co/Qwen/Qwen2.5-VL-7B-Instruct for experimental verification. During the SFT stage, we fine-tune both the visual projection layers and the LLM components by jointly predicting ThinkVG paths and task-related results. Particularly, the model is trained for 2 epochs on 8 NVIDIA H100 GPUs with a batch size of 256. The network is warmed up in the first 0.05 epochs with a linear learning rate from 3e-7 to 1e-4, which further decays by cosine schedule. The optimizer is AdamW. To accelerate training, we employ the DeepSpeed ZeRO-2 and the bf16 (Brain Floating Point) data format. During the RFT stage, the model is trained for 1 epoch on 8 NVIDIA H100 GPUs with a batch size of 128. The optimizer is AdamW with a learning rate of $62e-$ 6. Similar to SFT, we also fine-tune both the visual projector and LLM components. Regarding the hyper-parameters in Equation 2, we generate $8$ outputs (i.e., $G$) for each input and set $\beta$ and $\epsilon$ to $31e-$ 3 and $0.2$, respectively. The sample generator and LLM used for semantics reward are deployed by vLLM [14] framework.
5.3 Main Results
@@
We report the main results in Table 4. It can be seen that (1) Compared to zero-shot learning, fine-tuning can significantly improve the model’s performance; (2) When SFT is conducted with only 200K vanilla data (i.e., GEMeX-200K), the model lags behind the one trained on the full dataset by a large margin, with an average A-score that is around 4.5% lower; (3) When ThinkVG data (GEMeX-ThinkVG-200K) is further used for SFT, the model’s performance improves substantially, even surpassing that of the model trained on the full dataset in certain tasks, e.g., an improvement of approximately 5% on closed-ended tasks. Nevertheless, it can be observed that when only SFT is used, the improvement in visual grounding is quite limited. This is often due to inaccurate location reasoning in the thinking path after learning ThinkVG with SFT, which leads to error propagation in the final outputs; (4) To address this issue, further applying RFT with semantics reward and grounding reward yields noticeable enhancement, a 1.5%–3% improvement in both answer accuracy and localization performance. It is worth noting that, compared to using the full training data, the model further fine-tuned with RFT achieves better results in answer generation (i.e., higher average A-score) and comparable performance in localization. These findings demonstrate the effectiveness of our proposed ThinkVG dataset and RFT rewards.
@@
ThinkVG-200K 80.08 80.15 80.00 79.69 0.192
[…]
5.4 Robustness to Input Variations
We conduct two common perturbations to generate input variations: (1) we modify closed-ended questions while preserving their semantic integrity—for example, e.g., changing “Is the heart size abnormal in this CXR?” by randomly replacing the interrogative word “is” with “isn’t”, or replacing “abnormal” with “normal”, and adjusting the corresponding answer accordingly. During the test phase, each question has a 50% chance of being modified by one of the two modes mentioned above. If a question does not meet the criteria for either mode, no modification will be applied; (2) we randomly change the order of the options in single-choice questions. For each type of perturbation, we conduct three rounds of test and show models’ performance in Table 5. To more clearly compare the effectiveness of ThinkVG, we compare the performance of the models only after SFT. The results show that after fine-tuning the model with ThinkVG, its robustness against interference is significantly improved (i.e., lower standard deviation), highlighting another advantage.
5.5 Qualitative Evaluation
As shown in Figure 4 and 5, we compare the performance of models on several challenging examples, trained with and without ThinkVG on GEMeX-200K. When using ThinkVG, we present the model outputs after RFT stage. It can be seen that when using ThinkVG with proper tuning, not only are the answers more accurate, but the model also carefully thinks with analyzing corresponding regions in the image when addressing questions, leading to stronger overall interpretability.
@@ 5.5 Qualitative Evaluation
Figure 4: Challenging examples from GEMeX answered by models trained with and without ThinkVG. (✓) or (✗) in outputs highlight correct or incorrect reasons or answers. The colored words indicate the thinking with visual grounding process.
Figure 5: Challenging examples from GEMeX answered by models trained with and without ThinkVG. (✓) or (✗) in outputs highlight correct or incorrect reasons or answers. The colored words indicate the thinking with visual grounding process.
6 Conclusion
@@
In this work, we address the critical issue of interpretability in Medical Visual Question Answering by introducing a thinking with visual grounding dataset (ThinkVG). It explicitly links thinking steps to corresponding regions in the medical image, offering fine-grained visual explanations that enhance transparency and trust in model-generated answers. To further guide the model in generating accurate thinking paths, we proposed a new reward mechanism for reinforcement learning that aligns the model’s outputs with relevant visual and textual evidence. Despite using a subset of training data, the trained model is able to achieve impressive performance. We believe this work marks a step toward making Medical VQA systems more explainable and clinically usable, paving the way for safer and more reliable AI-assisted diagnosis.

### v2

@@ 4.2.1. Group Relative Policy Optimization
$\begin{aligned} \quad\quad\mathcal{J}_{\text{GRPO}}(\theta)&=\mathbb{E}_{q\sim P(Q),\,\{o_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}(O|q)}\Bigg[\\ \frac{1}{G}\sum_{i=1}^{G}\min\Bigg(\frac{\pi_{\theta}(o_{i}\mid q)}{\pi_{\theta_{\text{old}}}(o_{i}\mid q)}A_{i},&\text{clip}\left(\frac{\pi_{\theta}(o_{i}\mid q)}{\pi_{\theta_{\text{old}}}(o_{i}\mid q)},1-\varepsilon,1+\varepsilon\right)A_{i}\Bigg)\\ &-\beta D_{\text{KL}}(\pi_{\theta}\parallel\pi_{\text{ref}})\Bigg]\end{aligned}$
where $A_{i}=\frac{r_{i}-\text{mean}(\{r_{1},r_{2},\ldots,r_{G}\})}{\text{std}(\{r_{1},r_{2},\ldots,r_{G}\})}$ is the relative advantage over group rewards $\{r_{i}\}_{i}^{G}$, $D_{\text{KL}}$ is the KL divergence between $\pi_{\theta}$ and $\pi_{\text{ref}}$ (the model after SFT stage) to prevent catastrophic forgetting caused by over optimization, $\epsilon$ and $\beta$ are the PPO clipping hyper-parameter (Schulman et al., 2017) and the coefficient for the KL penalty, respectively.
messages = [“role”:“user”, “content”: f“‘We would like to request your feedback on the performance of two AI assistants in response to the user question displayed above. For your reference, the visual content in the image is represented with caption describing the same image. Please rate the accuracy (most important), relevance of their responses, considering both answer and reason (if any). Each assistant receives an overall score on a scale of 1 to 10, where a higher score indicates better overall performance. Please output both scores and your reaon in JSON format { assistant1: score, assistant2: score, reason:your_reason }.”””
Table 3. Our designed prompt for Semantics Reward.
@@
GEMeX-RMCoT-200K 94.71 56.52 87.11 66.23 80.08 61.76 70.20 51.13 83.03
GEMeX-RMCoT-200K 96.44 59.63 90.24 68.39 83.31 64.52 73.18 54.49 85.79
Table 4. Performance of Qwen2.5-VL-7B with different evaluating paradigms on GEMeX. The A-score indicates answer or choice accuracy (%), and the V-score represents mIoU (%). $\dagger$} indicates that the accuracy of the answer is judged by the LLM, as the question is open-ended. The best results are bolded, and the second-best are underlined in each column.
4.2.2. Reward Functions
We propose two reward functions, tailored for the two output parts of GEMeX: answer and involved visual locations. (1) Semantics Reward: In GEMeX, four types of questions are defined (Table 2). For choice-based and closed-ended questions, reward design is straightforward, as we can directly use Accuracy Reward (Pan et al., 2025) to evaluate whether the model’s outputs are correct. Nevertheless, designing a reward function for open-ended questions is particularly challenging due to the absence of fixed answers. Unlike (Liang, 2025) which employs traditional machine learning metrics such as BLEU as rewards—metrics that have been shown to poorly capture essential medical content (Jain et al., 2021)—we introduce a semantics-aware accuracy reward. It enables the integration of all four types into a unified, coherent reward paradigm. Specifically, we input both the ground truth (GT) answer and the model-generated answer into OpenBioLLM-70B, which can provide two accuracy scores at the semantics level. When these two scores are close, i.e., if the absolute difference between them is less than 2, we assign a reward of 1 point; otherwise, the reward is 0. The specific prompt is shown in Table 3. (2) Grounding Reward: Besides giving answers, the model needs to provide the visual locations involved in solving the questions. Therefore, we first check whether the number of bounding boxes predicted by the model matches the number in the GT. If they match, we then evaluate whether the mean Intersection over Union (mIoU) exceeds 0.75. A reward score of 1.0 is assigned only when both conditions are satisfied; otherwise, the reward is 0.
5. Experiments
5.1. Dataset
@@
We conduct experiments on the GEMeX dataset and randomly sample a portion of its training set to generate RMCoT data. We denote the version with the complete training set of 1.59M VQA triples as GEMeX-Full. The 200K subset that includes RMCoT data is denoted as GEMeX-RMCoT-200K, while the subset without RMCoT is referred to as GEMeX-200K. For evaluation, we use the original GEMeX test set, which consists of 300 images and 3,960 questions (1,144 open-ended questions, 543 closed-ended questions, 1,300 single-choice questions, and 973 multiple-choice questions).
5.2. Training Details
@@
We mainly explore the Qwen2.5-VL-7B-Instruct for experimental verification. During the SFT stage, we fine-tune both the visual projection layers and the LLM components. Particularly, the model is trained for 2 epochs on 8 NVIDIA H100 GPUs with a batch size of 256. The network is warmed up in the first 0.05 epochs with a linear learning rate from 3e-7 to 1e-4, which further decays by cosine schedule. The optimizer is AdamW. During the RFT stage, the model is trained for 1 epoch on 8 NVIDIA H100 GPUs with a batch size of 128. The learning rate is 2e62e-6. Regarding the hyper-parameters in Eq. 2, we generate $8$ rollouts (i.e., $G$) for each input and set $\beta$ and $\epsilon$ to 1e31e-3 and $0.2$, respectively. The sample generator and LLM used for semantics reward are deployed by vLLM (Kwon et al., 2023).
5.3. Main Results
@@
We report the main results in Table 4. It can be seen that (1) Compared to zero-shot learning, fine-tuning can significantly improve the model’s performance; (2) When SFT is conducted with only 200K vanilla data (i.e., GEMeX-200K), the model lags behind the one trained on the full dataset by a large margin, with an average A-score that is around 4.5% lower; (3) When RMCoT data (GEMeX-RMCoT-200K) is further used for SFT, the model’s performance improves substantially, even surpassing that of the model trained on the full dataset in certain tasks, e.g., an improvement of approximately 5% on closed-ended tasks. Nevertheless, it can be observed that when only SFT is used, the improvement in visual grounding is quite limited. This is often due to inaccurate location reasoning in the thinking path after learning RMCoT with SFT, which leads to error propagation in the final outputs; (4) To address this issue, further applying RFT with semantics reward and grounding reward yields noticeable enhancement, a 1.5%–3% improvement in both answer accuracy and localization performance. It is worth noting that, compared to using the full training data, the model further fine-tuned with RFT achieves better results in answer generation (i.e., higher average A-score) and comparable performance in localization. These findings demonstrate the effectiveness of our proposed RMCoT dataset and RFT rewards.
@@
RMCoT-200K 80.08 80.15 80.00 79.69 0.192
Table 5. Model performance under different perturbations. $\dagger$} means the standard deviation across the three perturbations from original A-score.
@@
Figure 4. One challenging example from GEMeX answered by models trained with and without RMCoT. (✓) or (✗) in outputs highlight correct or incorrect reasons or answers. The colored words indicate the thinking with visual grounding process.
5.4. Robustness to Input Variations
We conduct two common perturbations to generate input variations: (1) we modify closed-ended questions while preserving their semantic integrity—for example, changing “Is the heart size abnormal in this CXR?” by randomly replacing the interrogative word “is” with “isn’t”, or replacing “abnormal” with “normal”, and adjusting the corresponding answer accordingly. During the test phase, each question has a 50% chance of being modified by one of the two modes mentioned above. If a question does not meet the criteria for either mode, no modification will be applied; (2) we randomly change the order of the options in single-choice questions. For each type of perturbation, we conduct three rounds of test and show models’ performance in Table 5. To more clearly compare the effectiveness of RMCoT, we compare the performance of the models only after SFT. The results show that after fine-tuning the model with RMCoT, its robustness against interference is significantly improved (i.e., lower standard deviation), highlighting another advantage, especially in the context of potential challenges encountered in practical deployment.
5.5. Qualitative Evaluation
As shown in Figure 4, we compare the performance of models trained with and without RMCoT on GEMeX-200K. When using RMCoT, we present the model outputs after RFT stage. It can be seen that when using RMCoT with proper tuning, not only are the answers more accurate, but the model also carefully thinks with analyzing corresponding regions in the image when addressing questions, leading to stronger overall interpretability. More examples are shown on the project page.
6. Conclusion
@@
In this work, we address the critical issue of interpretability in Medical Visual Question Answering by introducing a region-aware multimodal chain-of-thought dataset (RMCoT). It explicitly links thinking steps to corresponding regions in the medical image, offering fine-grained visual explanations that enhance transparency and trust in model-generated answers. To further guide the model in generating accurate thinking paths, we proposed a new reward mechanism for reinforcement learning that aligns the model’s outputs with relevant visual and textual evidence. Despite using a subset of training data, the trained model is able to achieve impressive performance. We believe this work marks a step toward making Med-VQA systems more explainable and clinically usable, paving the way for safer and more reliable AI-assisted diagnosis.
Acknowledgements.
We thank all reviewers for their valuable time and feedback. This work was supported by Project P0056021 under Parent Project P0050643 of the Otto Poon Charitable Foundation Smart Cities Research Institute, HK PolyU; the Research Student Attachment Programme, HK PolyU; and the A*STAR Central Research Fund (“Robust and Trustworthy AI system for Multi-modality Healthcare”), Singapore.